Pixel:

Isabella is solving the equation 4x^2=13x−3 with the quadratic formula. Which values could she use for a, b, and c? a = 4, b = 3, c = −13 a = 4, b = −3, c = 13 a = 4, b = 13, c = −3 a = 4, b = −13, c = 3

Pixel:

@dude

dude:

We have to set the equation equal to 0

Pixel:

A den?

dude:

No lol

Pixel:

;-;

Pixel:

mk gimme a sec

Pixel:

AH C

dude:

Welp \(4x^2=13x−3\) You want to have the equation set to 0 Subtract 13x and add 3 on both sides! \(4x^2-13x+3=0\) This equation is written in \(ax^2+bx+c\) So... \(a=1,~b=-13,~c=3\)

Pixel:

but there is no a=1 .-..

dude:

\(a=4\) My mistake x'D

Pixel:

I SAID THAT YOU DING DONG

dude:

That is answer choice D.